1. (3x^2-2x+10)-(3x^2+3x+1)= 2. (y-y^2)-(2y-2y^2)-(3+y)= 3. (2w^3-3w^2+4)-(-2w^2+7w+1)= 4. (-x^7+10x+7)-(9x^7-4x-1)= 5. (x^4+3x^3-2x-19)-(-x^5+4x^4-3x-3)= 6. (2x^2+9x+1)-(x^2+4)= 7.(2a^3-3a^2+a)-(2a^3-3a^2+a+1)= 9.(11s-1)(11s+1)= 10.(8q+3pq)(8q-3pq)= 11.(mn+5n^2)(-5n^2+mn)= 12.(4xyz-2xy)(4xyz+2xy)= 13.(x-y)(x+y)(x^2+y^2)= 14.(4a-3b)(4a+3b)= 15.(kx+5)(kx-5)= 16.(2a^3-b^2)(2a^3+b^2)= 17.(3a+5x)(3a-5x)= 18.(mn-m^2)(-n+m)= 19.(-5k+3b)(b+2k)= 20.(-5k-3b)(-b-2k)= 21.(-2a+3b)(a^2b+b^2a)=
Answer by Alan3354(69443)

(3x^2-2x+10)-(3x^2+3x+1)= 2.
-5x + 9 = 2
-5x = -7
x = 1.4